# Optimas: Optimizing Compound AI Systems with Globally Aligned Local Rewards

### v2

@@ Appendix A Theoretical Analysis
According to the procedure described in Section 4, the positive and negative pairs are determined by comparing the expected task metrics. We assume the estimated metrics at $y_{k}^{+}$ and $y_{k}^{-}$ are chosen following (yk+ is labeled as $\mathbb{P}(\text{$y_{k}^{+}$ is labeled as positive})=\sigma_{\alpha}(\mathbb{E}_{\text{downstream}}\Bigl{[}R\bigl{(}x,f(x;\mathbf{v}_{-k}(x))\bigr{)}\,\big{|}\,y_{k}^{+}\Bigr{]}-\mathbb{E}_{\text{downstream}}\Bigl{[}R\bigl{(}x,f(x;\mathbf{v}_{-k}(x))\bigr{)}\,\big{|}\,y_{k}^{-}\Bigr{]})$, where $\sigma_{\alpha}(u)=\frac{1}{1+\exp(-\alpha u)}$ is the sigmoid function with parameter $\alpha>0$. $\alpha=+\infty$ corresponds to the case where the pairs are chosen deterministically.
Theorem A.1.
Under the conditions specified above, the maximizer of (5) satisfies the local-global alignment property (4). In addition, maximizing $r_{k}(x,C_{k}(x_{k};\mathbf{v}_{k}))$ over $v_{k}$ and maximizing $R(x,f(x;\mathbf{v}_{-k})\mid C_{k}(x_{k};\mathbf{v}_{k}))$ over $v_{k}$ will yield the same solution.
@@
We take the derivative of the left-hand side with respect to $p$ and set it to 00:
@@
Applying this lemma, we can obtain the solution to (5) is $\alpha\cdot\mathbb{E}_{\text{downstream}}\Bigl{[}R\bigl{(}x,f(x;\mathbf{v}_{-k}(x))\bigr{)}\,\big{|}\,y_{k}\Bigr{]}$ for some positive $\alpha$, and therefore it satisfies the local-global alignment property.(4)
@@
Lemma A.2.
Assume the local–global alignment property (4) holds for component $C_{k}$. Let $\mathbf{v}(x)$ and $\tilde{\mathbf{v}}(x)$ be two configuration policies that differ only in the policy for component $C_{k}$; denote the corresponding local outputs by $y_{k}$ and $\tilde{y}_{k}$, respectively. If
@@
where $\mathbf{v}_{-k}(x)$ denotes downstream configurations (independent of the choice at $C_{k}$). By assumption on the expected local reward, we have $r_{k}(x_{k},\tilde{y}_{k})\geq r_{k}(x_{k},y_{k})$ for almost every $x$. Applying the alignment property (4) pointwise yields
@@
Now when we consider all the iterations that are later than the time $T$, $\mathbf{v}^{(2)}$ is fixed, and we only need to consider the update regarding $\mathbf{v}^{(1)}$. In this case, we apply Theorem 4.1 of [58] and complete the proof.
Appendix B Dataset Details
This appendix provides the essential statistics and source information for the five compound-system benchmarks used in our experiments (Figure 2). For each dataset we specify the train / validation / test split sizes, the task formulation as used in the compound pipeline, and the evaluation metric reported in the main paper
Amazon (Behavior-Driven Next-Item Recommendation).
The corpus is derived from Amazon MMLU dataset Jin et al. [51]. Each instance consists of a user’s historical behaviour sequence (views, clicks, purchases) and the target “next item” to be recommended. We split it into to 335 / 60 / 99 user sequences after filtering malformed entries. Accuracy - whether the predicted item number 1 matches the ground truth - is the evaluation metric.
@@ Appendix B Dataset Details
PubMedQA [52] contains biomedical abstracts paired with yes/no/maybe answers to research questions. We keep the original “expert” split and discard ambiguous samples, resulting in 475 / 25 / 500 question–abstract pairs. Our compound system frames the task as three-way classification; exact-match accuracy is reported.
@@
STaRK-Prime -Prime originates from STARK benchmark introduced by Wu et al. [54]. It blends free-text passages with relational triples from biomedical knowledge graphs. Queries are natural-language questions; relevance labels are automatically propagated from the original STARK annotations. We uses the original dataset split: 495 / 51 / 96 queries. Performance is measured by Hit@1, which is the rate of ranking the ground truth items in the predicted ranking list.
@@
We adopt the HotpotQA [55] and keep the official train/dev/test splits: 1000, 250, and 100 questions respectively. Each example in the set contains a question and its (human-annotated) answer. We report answer-level F1 score.
@@
We use a subset of the full-instruction subset of BigCodeBench [56] due to efficiency issue. After proportionally drop the data, we obtain 500 / 25 / 70 coding tasks. Each sample includes a natural-language specification and reference unit tests. Our metric is pass@1: the proportion of generated programs that pass all tests in one try.
Appendix C Compound AI System Details
@@
[…]
@@
[…]
@@
Session Analyzer and Candidate Profiler both use the Qwen 2.5 1B model; we optimize their model parameters with PPO reinforcement learning [22]. This helps each module better encode task-specific knowledge, i.e., user sessions and product candidates. The final Next Item Decider is a GPT-4o-mini module, whose prompt we optimize.
@@
Appendix D Baseline Details
@@
  • LLMSelector: We run LLMSelector [8] with the LLMDiagnoser to estimate per-module performance, following their procedure. We perform two rounds of allocation updates with 100 training examples each round.
  • Hierarchical Behavior Cloning (HBC): We run HBC using the collected preference dataset by replacing the original reward model with the embedding similarity score. With the same input in the preference dataset, we use text-embedding-3-small to embed the module output and the preferred output in the preference dataset and calculate the embedding similarity score. We further weight the similarity score using the gap of the preferred output score and the rejected output score to get the reward for HBC.
  • TextGrad: We run TextGrad using GPT-4o mini to optimize each component’s prompt independently in separate epochs. Validation is performed every two optimization steps using 20 held-out validation instances. A batch size of 4 is used across all components and datasets. The best-performing prompt, as determined by validation frequency of LLM-generated textual feedback, is selected as the final configuration.
  • DSPy: We conduct optimization on DSPy’s MIPRO [23] prompt optimization approach. For fair comparison, we disable the few-shot example and system prompt optimization and only conducts optimization on the user instructions. We dynamically set the number of iterations for the MIPRO optimizer to match the budget of system runs in Table 3.
@@
Appendix E Optimas Details
We provide open access to our implementation code, including the Optimas framework and the compound system implementations for all five benchmarks at https://anonymous.4open.science/r/optimas-1308/.
Component selection. At each iteration $t$, we select a component to optimize. The selection can be made uniformly at random, or guided by a heuristic that prioritizes components likely to be performance bottlenecks. Specifically, for each component, we compute the average difference in expected global rewards between preferred and non-preferred outputs in its previous preference data. This value reflects the potential impact of improving the component, and we sample one for optimization with probability proportional to its average metric gap.
Local Optimization Steps for Different Configurations. Since each component controls only one configuration type, we can tailor the optimization approach as follows:
@@ Appendix E Optimas Details
  • Prompt tuning. For textual prompts, we apply prompt optimization algorithms [17, 19], using $r_{k}$ as the evaluation metric. We sample multiple prompts limited by a max number of prompt candidates. The prompts are ranked by average reward over validation instances, and the best-performing prompt is selected.
@@
  • Model fine-tuning. When $C_{k}$ is an LLM or neural model with trainable parameters, we can apply reinforcement learning algorithms, such as Proximal Policy Optimization (PPO) [22], using $r_{k}$ as the critic. The model parameters are updated for a small and fixed number of steps.
@@
  • Model selection and hyperparameter tuning. For discrete or low-dimensional continuous configurations, such as model selection, tool routing, or scalar hyperparameters, we formulate the optimization as a sampling problem parameterized by a probabilistic distribution. Since these configurations are instance-specific, the expectation in Eq. (7) reduces to a single input. For each input $x$, we evaluate a set of candidate configurations using the LRF $r_{k}$, and compute a probability distribution over candidates proportional to $\exp\{r_{k}(x_{k},C_{k}(x_{k};\mathbf{v}_{k}))\}$. This distribution is then used to sample the configuration update for the current iteration.
@@
Under a conservative update to the configuration of a component $C_{k}$, the expected global reward is guaranteed to maintain or improve, if the local–global alignment property in Eq. (4) holds.
Appendix F More Experiment Results
(a) Amazon system
(b) STaRK-Prime system
(c) HotpotQA system
Figure 7: Configuration updates during Optimas optimization on three compound AI systems: Amazon, HotpotQA, and STaRK-Prime .

### v3

@@ Appendix B Theoretical Analysis
According to the procedure described in Section 4, the positive and negative pairs are determined by comparing the expected task metrics. We assume the estimated metrics at $y_{k}^{+}$ and $y_{k}^{-}$ are chosen following $\mathbb{P}(\text{$y_{k}^{+}$ is labeled as positive})=\sigma_{\alpha}(\mathbb{E}_{\text{downstream}}\Bigl[R\bigl(x,f(x;\mathbf{v}_{-k}(x))\bigr)\,\big|\,y_{k}^{+}\Bigr]-\mathbb{E}_{\text{downstream}}\Bigl[R\bigl(x,f(x;\mathbf{v}_{-k}(x))\bigr)\,\big|\,y_{k}^{-}\Bigr])$, where $\sigma_{\alpha}(u)=\frac{1}{1+\exp(-\alpha u)}$ is the sigmoid function with parameter $\alpha>0$. $\alpha=+\infty$ corresponds to the case where the pairs are chosen deterministically.
Theorem B.1.
Under the conditions specified above, the maximizer of Eq.  4 satisfies the local-global alignment property Eq.  3. In addition, maximizing $r_{k}(x,C_{k}(x_{k};\mathbf{v}_{k}))$ over $v_{k}$ and maximizing $R(x,f(x;\mathbf{v}_{-k})\mid C_{k}(x_{k};\mathbf{v}_{k}))$ over $v_{k}$ will yield the same solution.
@@
We take the derivative of the left-hand side with respect to $p$ and set it to 0:
@@
Applying this lemma, we can obtain the solution to Eq.  4 is $\alpha\cdot\mathbb{E}_{\text{downstream}}\Bigl[R\bigl(x,f(x;\mathbf{v}_{-k}(x))\bigr)\,\big|\,y_{k}\Bigr]$ for some positive $\alpha$, and therefore it satisfies the local-global alignment property Eq.  3.
@@
Lemma B.2.
Assume the local–global alignment property Eq.  3 holds for component $C_{k}$. Let $\mathbf{v}(x)$ and $\tilde{\mathbf{v}}(x)$ be two configuration policies that differ only in the policy for component $C_{k}$; denote the corresponding local outputs by $y_{k}$ and $\tilde{y}_{k}$, respectively. If
@@
where $\mathbf{v}_{-k}(x)$ denotes downstream configurations (independent of the choice at $C_{k}$). By assumption on the expected local reward, we have $r_{k}(x_{k},\tilde{y}_{k})\geq r_{k}(x_{k},y_{k})$ for almost every $x$. Applying the alignment property Eq.  3 pointwise yields
@@
Now when we consider all the iterations that are later than the time $T$, $\mathbf{v}^{(2)}$ is fixed, and we only need to consider the update regarding $\mathbf{v}^{(1)}$. In this case, we apply Theorem 4.1 of (Tseng, 2001) and complete the proof.
Appendix C Dataset Details
This appendix provides the essential statistics and source information for the five compound‑system benchmarks used in our experiments (Figure 2). For each dataset we specify the train / validation / test split sizes, the task formulation as used in the compound pipeline, and the evaluation metric reported in the main paper
Amazon (Behavior‑Driven Next‑Item Recommendation).
The corpus is derived from Amazon MMLU dataset (Jin et al., 2024). Each instance consists of a user’s historical behaviour sequence (views, clicks, purchases) and the target “next item” to be recommended. We split it into to 335 / 60 / 99 user sequences after filtering malformed entries. Accuracy - whether the predicted item number 1 matches the ground truth - is the evaluation metric.
@@ Appendix C Dataset Details
PubMedQA (Jin et al., 2019) contains biomedical abstracts paired with yes/no/maybe answers to research questions. We keep the original “expert” split and discard ambiguous samples, resulting in 475 / 25 / 500 question–abstract pairs. Our compound system frames the task as three‑way classification; exact‑match accuracy is reported.
@@
STaRK-Prime ‑Prime originates from STARK benchmark introduced by (Wu et al., 2024b). It blends free‑text passages with relational triples from biomedical knowledge graphs. Queries are natural‑language questions; relevance labels are automatically propagated from the original STARK annotations. We uses the original dataset split: 495 / 51 / 96 queries. Performance is measured by Hit@1, which is the rate of ranking the ground truth items in the predicted ranking list.
@@
We adopt the HotpotQA (Yang et al., 2018) and keep the official train/dev/test splits: 1000, 250, and 100 questions respectively. Each example in the set contains a question and its (human-annotated) answer. We report answer‑level F1 score.
@@
We use a subset of the full-instruction subset of BigCodeBench (Zhuo et al., 2024) due to efficiency issue. After proportionally drop the data, we obtain 500 / 25 / 70 coding tasks. Each sample includes a natural‑language specification and reference unit tests. Our metric is pass@1: the proportion of generated programs that pass all tests in one try.
Appendix D Compound AI System Details
@@
Table 5: Modules, models, and optimization methods. Model Selection (LLMs) [^1] Model Selection (LLMs): we search over {gpt-4o, gpt-4o-mini, gpt-3.5-turbo-0125, gpt-4-turbo, claude-3-5-haiku-20241022, claude-3-5-sonnet-20241022, claude-3-7-sonnet-20250219}. indicates a discrete choice of LLMs. Aggregator [^2] Aggregator (STaRK): we search $\texttt{relation\_weight},\texttt{text\_weight}\in\{0.1,1.0\}$. is tuned over coefficients relation_weight and text_weight. Retriever [^3] Retriever (HotpotQA): we search $\texttt{k}\in\{1,5,10,25\}$. has a hyperparameter k.
@@
Aggregator [^2] Aggregator (STaRK): we search $\texttt{relation\_weight},\texttt{text\_weight}\in\{0.1,1.0\}$. Coefficients Hyperparam Search
@@
Session Analyzer and Candidate Profiler both use the Qwen 2.5 1B model; we optimize their model parameters with PPO reinforcement learning (Schulman et al., 2017). This helps each module better encode task-specific knowledge, i.e., user sessions and product candidates. The final Next Item Decider is a GPT-4o-mini module, whose prompt we optimize.
@@
Appendix E Baseline Details
@@
  • LLMSelector (Chen et al., 2025b): A lightweight policy selects the best LLM per component via model routing, without updating other configurations. Only applicable on PubMedQA.
    We run LLMSelector (Chen et al., 2025b) with the LLMDiagnoser to estimate per-module performance, following their procedure. We perform two rounds of allocation updates with 100 training examples each round.
  • Hierarchical Behavior Cloning (HBC) (Le et al., 2018): A hierarchical imitation learning method that optimizes components to produce outputs similar to those that lead to high global rewards.
    We run HBC using the collected preference dataset by replacing the original reward model with the embedding similarity score. With the same input in the preference dataset, we use text-embedding-3-small to embed the module output and the preferred output in the preference dataset and calculate the embedding similarity score. We further weight the similarity score using the gap of the preferred output score and the rejected output score to get the reward for HBC.
  • TextGrad (Yuksekgonul et al., 2025): A gradient-based prompt tuning method using estimated gradients from black-box LLMs to improve prompt efficacy.
    We run TextGrad using GPT-4o mini to optimize each component’s prompt independently in separate epochs. Validation is performed every two optimization steps using 20 held-out validation instances. A batch size of 4 is used across all components and datasets. The best-performing prompt, as determined by validation frequency of LLM-generated textual feedback, is selected as the final configuration.
  • DSPy (Khattab et al., 2023; Opsahl-Ong et al., 2024): A prompt optimization framework using the MIPRO algorithm that jointly refines module-level instructions and few-shot demonstrations. We conduct optimization on DSPy’s MIPRO (Opsahl-Ong et al., 2024) prompt optimization approach. For fair comparison, we disable the few-shot example and system prompt optimization and only conducts optimization on the user instructions. We dynamically set the number of iterations for the MIPRO optimizer to match the budget of system runs in Table 3.
@@
Appendix F Optimas Details
Component selection. At each iteration $t$, we randomly select a component to optimize.
Local Optimization Steps for Different Configurations. Given a globally aligned LRF $r_{k}$, we perform local optimization on each component $C_{k}$ to improve its configuration $\mathbf{v}_{k}$. Specifically, we solve:
In practice, explicitly setting the trust region threshold $\delta$ can be difficult due to heterogeneous configuration types (e.g., continuous weights or discrete tokens). Instead, we adopt a conservative number of update steps to restrict the magnitude of change during each iteration.
@@ Appendix F Optimas Details
  • Prompt tuning. For textual prompts, we apply prompt optimization algorithms (Yang et al., 2024; Wu et al., 2024a), using $r_{k}$ as the evaluation metric. We sample multiple prompts limited by a max number of prompt candidates. The prompts are ranked by average reward over validation instances, and the best-performing prompt is selected.
@@
  • Model fine-tuning. When $C_{k}$ is an LLM or neural model with trainable parameters, we can apply reinforcement learning algorithms, such as Proximal Policy Optimization (PPO) (Schulman et al., 2017), using $r_{k}$ as the critic. The model parameters are updated for a small and fixed number of steps.
@@
  • Model selection and hyperparameter tuning. For discrete or low-dimensional continuous configurations, such as model selection, tool routing, or scalar hyperparameters, we formulate the optimization as a sampling problem parameterized by a probabilistic distribution. Since these configurations are instance-specific, the expectation in Eq. Eq.  5 reduces to a single input. For each input $x$, we evaluate a set of candidate configurations using the LRF $r_{k}$, and compute a probability distribution over candidates proportional to $\exp\{r_{k}(x_{k},C_{k}(x_{k};\mathbf{v}_{k}))\}$. This distribution is then used to sample the configuration update for the current iteration.
@@
Under a conservative update to the configuration of a component $C_{k}$, the expected global reward is guaranteed to maintain or improve, if the local–global alignment property in Eq. Eq.  3 holds.
Appendix G More Experiment Results
G.1 Optimas Data efficiency
Figure 7: Impact of training data size on evaluation accuracy for the HotpotQA system. Optimas maintains robust performance even with reduced training data, demonstrating only a 5.19% absolute performance drop when using just 12.5% of the data compared to the full dataset.
To evaluate the data efficiency of Optimas, we examine how the amount of training data (system runs) affects the performance of learned LRFs. We train local reward models on the HotpotQA system using varying percentages of available training data: 12.5%, 25%, 50%, 75%, and 100%. Each model is trained for 3 epochs, and performance is measured by pairwise ranking accuracy—how often an output with higher global reward is assigned a higher local reward compared to an output with lower global reward. Optimas maintains reasonable performance across different data sizes as show in Figure 7. Even with only 12.5% of the data, the system achieves 65.46% accuracy (92.7% of full-data performance), while 25% of the data yields 67.22% accuracy (95.1% of full-data performance). The diminishing returns beyond 25% data suggest that the core patterns for local-global alignment can be captured efficiently from limited preference data.
This data efficiency is particularly valuable in practice, where collecting extensive preference data through system runs can be computationally expensive. Combined with the reduced computational cost of local optimization, these results demonstrate that Optimas provides a practical solution for optimizing complex compound AI systems even in resource-constrained environments.
G.2 Local vs. Global reward landscapes
Table 8: On HotpotQA, sweeping the retriever’s top-$k$ reveals closely aligned local and global reward landscapes.
$k$ 1 2 3 5 10 15 25
Local reward 0.4247 0.5578 0.5695 0.6124 0.6117 0.5949 0.5123
Global reward 0.3398 0.3493 0.3325 0.3598 0.3645 0.3568 0.3465
To better understand how local objectives reflect global performance, we sweep the retriever’s top-$k$ setting on HotpotQA and compare local and global rewards (Table 8). The two landscapes are closely aligned: both are unimodal and peak at nearby values (k=5k{=}5 for the local reward and k=10k{=}10 for the global reward). The top-3 configurations $\{5,10,15\}$ coincide, differing only in the order of the top-2. This alignment shows that the local reward provides a reliable proxy for the global objective, lending empirical support to our theoretical guarantee that local optimization drives system-level gains.
G.3 LRF backbone size
Table 9: On HotpotQA, effect of LRF backbone size on pairwise ranking accuracy (%). Lightweight backbones are competitive with larger ones.
Backbone size Pairwise accuracy (%)
1B 72.58
3B 70.04
8B 71.20
We also examine whether the size of the LRF backbone affects alignment quality. On HotpotQA, we compare 1B, 3B, and 8B backbones (Table 9). Performance remains comparable across sizes, with pairwise accuracies of 72.58%, 70.04%, and 71.20%, respectively. The lack of clear benefit from larger backbones suggests that lightweight models are sufficient for this task, making local optimization both effective and computationally efficient.
G.4 Prompt candidates per step
Table 10: On HotpotQA, we vary the number of candidate prompts per optimization step. We compute the global performance (F1) under each experiments.
# candidates 3 5 7 10
Final F1 0.2822 0.1945 0.2968 0.2405
Another factor is the number of candidate prompts considered at each optimization step. On HotpotQA, we vary this number from 3 to 10 and measure the resulting global performance (Table 10). The best F1 score is achieved with 7 candidates, although results with fewer candidates remain competitive. These findings indicate that exhaustive candidate pools are unnecessary, and that modest numbers already yield strong performance with lower computational cost.
G.5 New inputs for LRF adaptation
Table 11: On HotpotQA, number of new inputs used to collect preference pairs for LRF adaptation vs. final global performance (F1).
# new inputs 10 20 30 40
Final F1 0.2773 0.2822 0.2659 0.2533
Finally, we explore how many new inputs are needed when adapting the LRFs. On HotpotQA, we test adaptation with 10 to 40 new inputs (Table 11). Performance improves up to about 20 inputs, after which gains plateau and even slightly decline. This suggests that effective adaptation can be achieved with relatively small amounts of new data, reinforcing the practicality of our approach in scenarios where data collection is limited.